# 1]Independent Researcher

### v1

@@ 3.1.2 Discovery of two-soliton solutions
[…]
@@
Here, $\eta$, $\sigma_{1}$, $\sigma_{2}$, $\rho_{1}$, $\rho_{2}$, $\hat{A}$, $\hat{B}$, and $\hat{C}$ are all real constants, $\mu$ and $\lambda$ are complex constants, and $\theta_{1}$ and $\theta_{2}$ are the real ones and can be varied within the range $[0,2\pi]$. $w_{1}(t)$ and $w_{2}(t)$ are real functions of the temporal variable $t$. It has been investigated that the amplitudes of the two circularly polarized waves are affected by $\mu$ and $\lambda$.
  • Parallel dark two-solitons
    By setting the perturbation coefficient $\alpha(t)=\frac{1}{2}t$, and choosing the other constants as $\eta=1,\sigma_{1}=1,k=4,\mu=\lambda=1+i,\theta_{1}=\frac{\pi}{4}$, and $\theta_{2}=\frac{\pi}{3}$ in (17)-(23), one can obtain the parallel dark two-soliton solutions to the (2+ 1)-dimensional VC-CNLSE (1). The nonlinearity coefficient can be considered proportional to the dispersion coefficient, expressed as $\beta(t)=-\eta\alpha(t)$. For data-driven dark two-soliton solutions to the model equation (1), we consider the spatial and temporal domains as $\Omega=[-8,8]\times[-6,6]$ and $[-0.1,0.1]$, respectively. In order to generate the initial and boundary training dataset, we consider the following equations
    $\begin{cases}\psi_{j}(x,y,t=-0.1)=\psi_{j}^{ic}(x,y,t=-0.1),\quad(x,y)\in\Omega,\\ \psi_{j}(x=-8,y,t)=\psi_{j}^{bc}(x=-8,y,t),\quad(y,t)\in[-6,6]\times(-0.1,0.1],\\ \psi_{j}(x,y=-6,t)=\psi_{j}^{bc}(x,y=-6,t),\quad(x,t)\in[-8,8]\times(-0.1,0.1],\\ \psi_{j}(x=8,y,t)=\psi_{j}^{bc}(x=8,y,t),\quad(y,t)\in[-6,6]\times(-0.1,0.1],\\ \psi_{j}(x,y=6,t)=\psi_{j}^{bc}(x,y=6,t),\quad(x,t)\in[-8,8]\times(-0.1,0.1],\quad j=1,2,\end{cases}$
@@
    where $\psi_{j}^{ic}$ and $\psi_{j}^{bc}$ can be constructed using the solutions provided in (17) and (18) for $j=1,2$. First of all, we divide the computational domain $\Omega\times[-0.1,0.1]$ into $256\times 256\times 201$ discrete equidistant points. We randomly generate $N_{ic}=200$ initial points and $N_{bc}=200$ boundary points, and use them to obtain the data-driven solutions through the use of the initial and boundary conditions described above. We also generate $N_{r}=10,000$ residual points by adopting the same sampling strategy used for the previous examples. The RAR strategy is used as previously mentioned. We aim to recover the dark two-soliton solutions given by (17) and (18) by applying the enhanced PINN framework. In doing so, we use a neural network model with $4$ hidden layers and $100$ neurons per layer. The underlying PINN model aims to learn the hidden patterns from the training data by minimizing the loss functions corresponding to the initial and boundary conditions, the residuals of the governing equations, and the slope recovery term. After $30,000$ iterations of the Adam optimizer with learning rate $0.0001$ over $2132.0729\,s$, we continue the optimization with $5,215$ iterations of the L-BFGS optimizer over $878.2534\,s$, and precisely simulate the dark two-soliton solutions to the system of equations (1). During this process, our model successfully reduced the $\mathbb{L}^{2}$ relative errors for $\psi_{1}$ and $\psi_{2}$ from $0.31\%$ and $0.32\%$ to $0.22\%$ and $0.25\%$, respectively. In Fig. 9, the density plots of the exact and predicted dynamics in the $x-t$ and $y-t$ planes (top row) and the $x-y$ plane (middle row) are displayed, showing that the parallel dark two-solitons do not interact with each other. Additionally, we compare the exact and predicted solutions at different temporal levels in the bottom row of Fig. 9, which shows that our trained model successfully captured the data-driven solutions to the system of equations (1). The evolution of the predicted dynamics in the $x-t$ plane (for $y=6$) and the $y-t$ plane (for $x=8$), along with the corresponding point-wise errors, are displayed through three-dimensional plots in Fig. 10. The behavior of the loss functions during training with the Adam optimizer and the L-BFGS optimizer is illustrated in Fig. 11. In the left panel, the loss curves corresponding to the overall loss function, initial condition loss function, boundary condition loss function, and residual loss function are displayed, showing a slow decrease in loss values with many oscillations. The final loss values for each loss function after training with the Adam optimizer are recorded, respectivelly, as follows: $022.751e-$ 02, $032.262e-$ 03, $048.808e-$ 04, and $021.235e-$ 02. From the middle panel, one can see that the individual residual losses behave relatively close to each other. In the right panel of Fig. 11, the L-BFGS optimizer exhibits smooth behavior of the loss values during training, with the final loss values recorded as follows: $038.103e-$ 03 for the overall loss function, $052.397e-$ 05 for the initial condition loss, $058.341e-$ 05 the boundary condition loss, and $034.409e-$ 03 for the residual loss.
@@
    Figure 10: Three-dimensional plots with contour maps and the corresponding point-wise errors for $\psi_{j}$ in the $x-t$ plane (first and second panels) and $y-t$ plane (third and fourth panels), from left to right.
@@
  • Parallel cubic dark two-solitons
    By taking $\eta=1,\sigma_{1}=1,k=4,\mu=\lambda=1+i,\theta_{1}=\frac{\pi}{4},\theta_{2}=\frac{\pi}{3}$, and choosing the dispersion coefficient as $\alpha(t)=\frac{1}{5}t^{2}$, the cubic dark two-dolitons to the (2+ 1)-dimensional VC-CNLSE (1) can be obtained. Here, the nonlinearity coefficient is considered as $\beta(t)=-\eta\alpha(t)$. We consider the following initial and boundary conditions for this type of soliton dynamics, with the spatial and temporal domains defined as $\Omega=[-10,10]\times[-6,6]$ and $[-2,2]$, respectively. Then, for $j=1,2$ we have
    $\begin{cases}\psi_{j}(x,y,t=-2)=\psi_{j}^{ic}(x,y,t=-2),\quad(x,y)\in\Omega,\\ \psi_{j}(x=-10,y,t)=\psi_{j}^{bc}(x=-10,y,t),\quad(y,t)\in[-6,6]\times(-2,2],\\ \psi_{j}(x,y=-6,t)=\psi_{j}^{bc}(x,y=-6,t),\quad(x,t)\in[-10,10]\times(-2,2],\\ \psi_{j}(x=10,y,t)=\psi_{j}^{bc}(x=10,y,t),\quad(y,t)\in[-6,6]\times(-2,2],\\ \psi_{j}(x,y=6,t)=\psi_{j}^{bc}(x,y=6,t),\quad(x,t)\in[-10,10]\times(-2,2],\end{cases}$
@@
    where $\psi_{j}^{ic}$ and $\psi_{j}^{bc}$ can be constructed, considering the provided constant values and the dispersion and nonlinearity coefficients for the dark two-soliton solutions (17)-(23). As before, we first divide the computational domain $\Omega\times[-2,2]$ into $256\times 256$ discrete equidistant points. We set $N_{ic}=200$ and $N_{bc}=200$ corresponding to the initial and boundary training data to be randomly chosen. In order to embed the physical laws described by the governing equations into the model’s knowledge, we generate $N_{r}=10,000$ residual points using the LHS strategy within the computational domain. We trained a PINN model based on the enhanced framework, equipped with a neural network model with the same structure as in the previous example. We used $30,000$ iterations of the Adam optimizer and $5,034$ iterations of the L-BFGS optimizer, with training times of $2135.0391\,s$ and $870.8201\,s$, respectively. Our trained model achieved the $\mathbb{L}^{2}$ relative errors of $0.64\%$ and $0.65\%$ for $\psi_{1}$ and $\psi_{2}$, respectively, with the Adam optimizer, and $0.15\%$ and $0.16\%$ with the L-BFGS optimizer. The dynamical behaviors of the exact and predicted cubic dark two-solitons are illustrated in the $x-t$ and $y-t$ planes in the top row, and the $x-y$ plane in the middle row of Fig. 12. One can see that the dark two-solitons do not interact with each other during propagation. To demonstrate that our trained model accurately captures the exact dynamics, comparisons between the predicted dynamics at different temporal levels are presented in the bottom row of Fig. 12. In Fig. 13, the evolution of the predicted dynamics in the $x-t$ plane (for $y=6$) and the $y-t$ plane (for $x=10$), with the corresponding point-wise errors, is displayed through three-dimensional plots. The loss values during the training phase with the Adam optimizer and L-BFGS optimizer are illustrated in Fig. 14. It can be seen that, the Adam optimizer exhibits oscillatory behavior in the loss functions during training, whereas the L-BFGS optimizer shows gradual descent without such oscillations. The loss values for the overall loss function, initial condition loss function, boundary condition loss function, and residual loss function after training with the Adam optimizer are recorded as follows: $021.608e-$ 02, $057.672e-$ 05, $059.445e-$ 05, and $048.703e-$ 04, respectively. Continuing the optimization process with the L-BFGS optimizer reduced these loss values to: $036.504e-$ 03, $062.494e-$ 06, $064.660e-$ 06, and $055.238e-$ 05, respectively. In the middle panel of Fig. 14, the behavior of the individual residual losses is depicted, showing that these loss values are relatively close to one another.
@@
    Figure 13: Three-dimensional plots with contour maps and the corresponding point-wise errors for $\psi_{j}$ in the $x-t$ plane (first and second panels) and $y-t$ plane (third and fourth panels), from left to right.
@@ 3.2 Data-driven parameter discovery to (2+ 1)-dimensional VC-CNLSE
  • Data-driven constant parameter discovery
    For the constant parameter discovery, we use the introduced enhanced PINN framework to predict the constant coefficient $\tau$ through training a PINN model. Given the same parameters for the parabolic vector dark one-soliton, we discretize the spatial and temporal domains into $256\times 256\times 201$ equidistant points. We randomly sample $N_{o}=1,000$ points from the generated data, then create a set of observations of the exact solutions $\Big{\{}u_{1}^{k},v_{1}^{k},u_{2}^{k},v_{2}^{k}\Big{\}}_{k=1}^{N_{o}}$ based on the sampled data points. The data-driven parameter discovery to predict the constant coefficient $\tau$ will be performed by training an enhanced PINN model, which is equipped with a neural network containing $5$ hidden layers with $40$ neurons per layer, through an optimization process that minimizes the following loss function:
@@
    $+\frac{1}{N_{r}}\sum_{j=1}^{N_{r}}w_{r}^{(k)}\Big{(}\Big{\lvert}\mathcal{F}_{\hat{u}_{1,\boldsymbol{\theta}}}(\mathbb{X}^{j},t^{j})\Big{\rvert}^{2}+\Big{\lvert}\mathcal{F}_{\hat{v}_{1,\boldsymbol{\theta}}}(\mathbb{X}^{j},t^{j})\Big{\rvert}^{2}$
    $+\Big{\lvert}\mathcal{G}_{\hat{u}_{2,\boldsymbol{\theta}}}(\mathbb{X}^{j},t^{j})\Big{\rvert}^{2}+\Big{\lvert}\mathcal{G}_{\hat{v}_{2,\boldsymbol{\theta}}}(\mathbb{X}^{j},t^{j})\Big{\rvert}^{2}\Big{)}+\frac{w_{a}}{\frac{1}{D-1}\sum_{\ell=1}^{D-1}\exp(\frac{\sum_{k=1}^{N^{\ell}}a_{k}^{\ell}}{N^{\ell}})},$
@@
    where $\boldsymbol{\theta}$ represents the set of all trainable parameters, including the weights, biases, scalable parameters, and the parameter $\hat{\tau}$, which serves as a prediction of the true value of $\tau$. We initialized this trainable parameter as $\hat{\tau}=0$. Minimizing this loss function allows our PINN model to align its predictions with data-driven observations while embedding the physical laws into the model’s knowledge. Incorporating these physical laws into the training phase is the key feature of the PINN model to predict the true value of the constant coefficient $\tau$. We allow our model to learn the true value of the constant coefficient $\tau$ by treating this parameter as one of the trainable parameters within the underlying neural network model. We used $10,000$ iterations of the Adam optimizer with the default learning rate, followed by a run of the L-BFGS optimizer, to obtain the optimal values of the trainable parameters. During this process, the $\mathbb{L}^{2}$ relative error between the true value of $\tau$ and its prediction $\hat{\tau}$ decreased from $0.61\%$ to $0.013\%$, demonstrating that our trained model can predict the true value of $\tau$ with high accuracy. To evaluate the performance of the enhanced PINN model for data-driven parameter discovery in the (2+ 1)-dimensional VC-CNLSE under noisy conditions, we introduce different noise intensities to the observational data. For a fair comparison, we use the same initialization for all networks trained with different noise intensities. As reported in Table 1, the enhanced PINN model accurately predicts the true value of the constant coefficient $\tau$ even when the observational data is corrupted with noise. It is evident that increasing noise levels leads to higher prediction errors and negatively impacts loss convergence. The left panel of Fig. 15 depicts the process of our PINN model learning the true value of the constant coefficient $\tau$. During the iterations of the Adam optimizer (solid lines), followed by the L-BFGS optimizer (dashed lines), our PINN model gradually predicts the true value of $\tau$ with high accuracy. One can see that increasing the noise levels will have a negative impact on the predictions. Training under different noise intensities alters the loss values, as shown in the middle panel of Fig. 15. The $\mathbb{L}^{2}$ relative errors between the constant coefficient $\tau$ and its prediction $\hat{\tau}$ for different levels of added noise are illustrated in the right panel of Fig. 15. It can be seen that as the noise level increases from 00 to $15$, the relative error also increases, and continuing the optimization with the L-BFGS optimizer does not have a significant effect on prediction accuracy. In Tabel 2, we investigate the effect of the number of observations on the prediction accuracy through the lens of $\mathbb{L}^{2}$ relative error. It can be seen that increasing the number of observations reduces the relative error, even in the presence of noisy data. In summary, our enhanced PINN model achieves lower prediction error and faster convergence with clean data (noise $=0\%$). However, introducing noise into the data negatively impacts the model’s prediction accuracy and convergence.
@@
[…]
    00 $0.9998723$ $0.013$ $038.536e-$ 03
    $5$ $0.9984364$ $0.16$ $021.934e-$ 02
    $10$ $0.98912466$ $1.09$ $024.896e-$ 02
    $15$ $1.01243$ $1.24$ $011.042e-$ 01
@@
  • Data-driven variable parameter discovery
@@
    Since these coefficients are functions of the temporal variable $t$, the corresponding neural network should process the temporal coordinate as input data to learn the dynamic of these coefficients. We divide the computational domain $\Omega\times[-2,2]$ into $256\times 256\times 200$ discrete equidistant points and randomly sample $N_{o}=2,000$ data observations from the generated data points. We also sample $N_{r}=20,000$ residual points using the LHS strategy within the computational domain. Additionally, we divide the temporal domain $[-2,2]$ into $1,000$ equidistant points to assess the accuracy of the underlying PINN model in predicting the variable coefficients through the lens of the $\mathbb{L}^{2}$ relative error. Our PINN model will be equipped with a neural network parameterized by $\boldsymbol{\theta}_{1}$, using spatial and temporal coordinates as input data to train and predict the solutions $\psi_{1}$ and $\psi_{2}$, along with another network parameterized by $\boldsymbol{\theta}_{2}$, which takes only the temporal coordinate as input data to predict the variable coefficients. Both networks may or may not have the same hidden layer structure and will be trained simultaneously. In this experiment, the first network has $5$ hidden layers with $40$ neurons in each, while the second network has $3$ hidden layers with $30$ neurons per layer. We consider the locally adaptive activation $\tanh$ function for both networks to accelerate the convergence speed. Then, the underlying PINN model aims to learn the dynamic behavior of both the solutions $\psi_{1}$ and $\psi_{2}$, as well as the dynamic behavior of the variable coefficients, through a minimization problem by minimizing the following loss function:

### v2

@@ 3.2.1 Discovery of two-soliton solutions
$k$ $:=\eta(\lvert\mu\rvert^{2}+\lvert\lambda\rvert^{2}),\quad\sigma_{2}:=\frac{\lvert\sin(\theta_{2})\rvert}{\lvert\sin(\theta_{1})\rvert}\sigma_{1},\quad b(t):=-k\int\alpha(t)dt,$ (15)
@@
Here, $\eta$, $\sigma_{1}$, $\sigma_{2}$, $\rho_{1}$, $\rho_{2}$, $\hat{A}$, $\hat{B}$, and $\hat{C}$ are all real constants, $\mu$ and $\lambda$ are complex constants, and $\theta_{1}$ and $\theta_{2}$ are the real ones and can be varied within the range $0,2\pi]$. $w_{1}(t)$ and $w_{2}(t)$ are real functions of the temporal variable $t$. It has been investigated that the amplitudes of the two circularly polarized waves are affected by $\mu$ and $\lambda$.
Case 3: Parallel dark two-solitons
@@
By setting the perturbation coefficient $\alpha(t)=\frac{1}{2}t$, and choosing the other constants as $\eta=1,\sigma_{1}=1,k=4,\mu=\lambda=1+i,\theta_{1}=\frac{\pi}{4}$, and $\theta_{2}=\frac{\pi}{3}$ in (13)-(19), one can obtain the parallel dark two-soliton solutions to the (2+ 1)-dimensional VC-CNLSE (1). The nonlinearity coefficient can be considered proportional to the dispersion coefficient, expressed as $\beta(t)=-\eta\alpha(t)$. The spatial and temporal domains are set as $\Omega=[-8,8]\times[-6,6]$ and [0.1,0.1][-0.1,0.1], discretized into $256\times 256\times 201$ equidistant points. We set $N_{ic}=200$ and $N_{bc}=200$, and generate $N_{r}=10,000$ residual points randomly using the LHS strategy. We aim to recover the dark two-soliton solutions given by (13) and (14) by applying the enhanced PINN framework. In doing so, we use a neural network model with $4$ hidden layers and $100$ neurons per layer. After $30,000$ iterations of the Adam optimizer with learning rate $0.0001$, followed by $5,215$ iterations of the L-BFGS optimizer the $\mathbb{L}^{2}$ relative errors for $\psi_{1}$ and $\psi_{2}$ are reduced from $0.31\%$ and $0.32\%$ to $0.22\%$ and $0.25\%$, respectively. In Fig. 11, the density plots of the exact and predicted dynamics in the xtx-t and yty-t planes (top row) and the xyx-y plane (middle row) are displayed, confirming that the parallel dark two-solitons propagate without interaction. Additionally, we compare the exact and predicted solutions at different temporal levels in the bottom row of Fig. 11, which shows that our trained model successfully captured the data-driven solutions to the system of equations (1). The evolution of the predicted dynamics in the xtx-t plane (for $y=6$) and the yty-t plane (for $x=8$), along with the corresponding point-wise errors, are displayed in Fig. 12. The behavior of the loss functions during training with the Adam optimizer and the L-BFGS optimizer is illustrated in Fig. 13. In the left panel, the loss trajectories through the optimization process using the Adam optimizer are illustrated, showing a slow decrease in loss values with many oscillations. From the middle panel, one can see that the individual residual losses behave relatively close to each other. In the right panel of Fig. 13, the L-BFGS optimizer exhibits smooth behavior of the loss values during training. In Table 4, a comparison between the proposed enhanced PINN and the method in [50] and the standard PINN is presented. All networks have the same structure, and $N_{r}=16,200$ is set for the two other methods. Additionally, the method in [50] uses the same activation functions as the enhanced PINN. One can see that, the enhanced PINN achieves the lowest $\mathbb{L}^{2}$ relative errors ($0.22\%$ for $\psi_{1}$ and $0.25\%$ for $\psi_{2}$) and the lowest loss values, with a training time ($3010$ s) shorter than the method in [50] $(3393$ s) and comparable to the standard PINN. In contrast, both other methods show significantly higher errors and losses.
@@
Figure 12: Three-dimensional plots with contour maps and the corresponding point-wise errors for $\psi_{j}$ in the xtx-t plane (first and second panels) and yty-t plane (third and fourth panels), from left to right.
@@
Table 4: Comparison results in terms of the $\mathbb{L}^{2}$ relative error, loss value, and training time (s).
$\mathbb{L}^{2}$ error Loss value
Method $\psi_{1}$ $\psi_{2}$ Adam L-BFGS Training time (s)
Enhanced PINN $0.22\%$ $0.25\%$ 2.7511e022.7511e-02 8.1037e038.1037e-03 $3010.3263$
Method in [50] $10.51\%$ $12.15\%$ 9.6591e009.6591e-$0$ 1.7260e001.7260e-$0$ $3392.7019$
PINN $17.62\%$ $19.03\%$ $15.0093e+01$ 9.1101e009.1101e-$0$ $3011.9031$
Case 4: Parallel cubic dark two-solitons
@@
By taking $\eta=1,\sigma_{1}=1,k=4,\mu=\lambda=1+i,\theta_{1}=\frac{\pi}{4},\theta_{2}=\frac{\pi}{3}$, and choosing the dispersion coefficient as $\alpha(t)=\frac{1}{5}t^{2}$, the cubic dark two-dolitons to the (2+ 1)-dimensional VC-CNLSE (1) can be obtained. Here, the nonlinearity coefficient is considered as $\beta(t)=-\eta\alpha(t)$. As before, we first divide the computational domain $10,10]\times[-6,6]\times[-2,2]$ into $256\times 256$ discrete equidistant points. We set $N_{ic}=200$ and $N_{bc}=200$ to be randomly chosen. We use the LHS strategy to generate $N_{r}=10,000$ residual points within the computational domain. Our enhanced PINN model comprises a network with $4$ hidden layers and $100$ neurons in each layer. After $30,000$ iterations of the Adam optimizer, followed by $5,034$ iterations of the L-BFGS optimizer the $\mathbb{L}^{2}$ relative errors for $\psi_{1}$ and $\psi_{2}$ are measured from $0.64\%$ and $0.65\%$ to $0.15\%$ and $0.16\%$, respectively. The dynamical behaviors of the exact and predicted cubic dark two-solitons are illustrated in the xtx-t and yty-t planes in the top row, and the xyx-y plane in the middle row of Fig. 14. One can see that the dark two-solitons do not interact with each other during propagation. Comparisons between predicted dynamics at different temporal levels are presented in the bottom row of Fig. 14. In Fig. 15, the evolution of the predicted dynamics in the xtx-t plane (for $y=6$) and the yty-t plane (for $x=10$), with the corresponding point-wise errors, is displayed through three-dimensional plots. The loss values during the training phase with the Adam optimizer and L-BFGS optimizer are illustrated in Fig. 16. It can be seen that, the Adam optimizer exhibits oscillatory behavior in the loss functions during training, whereas the L-BFGS optimizer shows gradual descent without such oscillations. In the middle panel of Fig. 16, the behavior of the individual residual losses is depicted, showing that these loss values are relatively close to one another. In Table 5, the performance of the proposed enhanced PINN is compared with the method in [50] and the standard PINN. All networks have the same structure, and $N_{r}=16,200$ is set for the other two methods. Additionally, the method in [50] uses the same activation functions as the enhanced PINN. The enhanced PINN achieves the lowest $\mathbb{L}^{2}$ relative errors and the lowest loss values for both Adam and L-BFGS optimizers. Its training time is also shorter than that of the method in [50] and similar to the standard PINN. In contrast, both the method in [50] and the standard PINN result in significantly higher errors and loss values.
@@
Figure 15: Three-dimensional plots with contour maps and the corresponding point-wise errors for $\psi_{j}$ in the xtx-t plane (first and second panels) and yty-t plane (third and fourth panels), from left to right.
@@
Table 5: Comparison results in terms of the $\mathbb{L}^{2}$ relative error, loss value, and training time (s).
$\mathbb{L}^{2}$ error Loss value
Method $\psi_{1}$ $\psi_{2}$ Adam L-BFGS Training time (s)
Enhanced PINN $0.15\%$ $0.16\%$ 1.6082e021.6082e-02 6.5041e036.5041e-03 $3005.8592$
Method in [50] $9.62\%$ $9.61\%$ 4.27301e004.27301e-$0$ 5.0026e015.0026e-01 $3209.9162$
PINN $13.04\%$ $13.05\%$ $5.28301e+01$ 8.2830e018.2830e-01 $3005.7590$
@@ 3.3 Data-driven parameter discovery to (2+ 1)-dimensional VC-CNLSE
$\bullet$ Data-driven constant parameter discovery
For the constant parameter discovery, we use the introduced enhanced PINN framework to predict the constant coefficient $\tau$ through training a PINN model. Given the same parameters for the parabolic vector dark one-soliton, we discretize the spatial and temporal domains into $256\times 256\times 201$ equidistant points. We randomly sample $N_{o}=1,000$ points from the generated data, then create a set of observations of the exact solutions $\Big{\{}u_{1}^{k},v_{1}^{k},u_{2}^{k},v_{2}^{k}\Big{\}}_{k=1}^{N_{o}}$ based on the sampled data points. We sample $N_{r}=10,000$ initial residual points. The data-driven parameter discovery to predict the constant coefficient $\tau$ will be performed by training an enhanced PINN model, which is equipped with a neural network containing $5$ hidden layers with $40$ neurons per layer, through an optimization process that minimizes the following loss function:
@@
$+\frac{1}{N_{r}}\sum_{k=1}^{N_{r}}\Big{(}w_{r}^{\mathcal{F}_{\hat{u}}}\Big{\lvert}\mathcal{F}_{\hat{u}_{1,\boldsymbol{\theta}}}(\mathbb{x}^{k},t^{k})\Big{\rvert}^{2}+w_{r}^{\mathcal{F}_{\hat{v}}}\Big{\lvert}\mathcal{F}_{\hat{v}_{1,\boldsymbol{\theta}}}(\mathbb{x}^{k},t^{k})\Big{\rvert}^{2}$
@@
$+w_{r}^{\mathcal{G}_{\hat{u}}}\Big{\lvert}\mathcal{G}_{\hat{u}_{2,\boldsymbol{\theta}}}(\mathbb{x}^{k},t^{k})\Big{\rvert}^{2}+w_{r}^{\mathcal{G}_{\hat{v}}}\Big{\lvert}\mathcal{G}_{\hat{v}_{2,\boldsymbol{\theta}}}(\mathbb{x}^{k},t^{k})\Big{\rvert}^{2}\Big{)}+w_{a}\cdot Loss_{a}(\boldsymbol{\theta}),$
@@
where $\boldsymbol{\theta}$ represents the set of all trainable parameters, including the weights, biases, scalable parameters, and the parameter $\hat{\tau}$, which serves as a prediction of the true value of $\tau$. We initialized this trainable parameter as $\hat{\tau}=0$. Minimizing this loss function allows our PINN model to align its predictions with data-driven observations while embedding the physical laws into the model’s knowledge. We allow our model to learn the true value of the constant coefficient $\tau$ by treating this parameter as one of the trainable parameters within the underlying neural network model. We used $10,000$ iterations of the Adam optimizer with the default learning rate, followed by a run of the L-BFGS optimizer, to obtain the optimal values of the trainable parameters. During this process, the $\mathbb{L}^{2}$ relative error between the true value of $\tau$ and its prediction $\hat{\tau}$ decreased from $0.61\%$ to $0.013\%$, demonstrating that our trained model can predict the true value of $\tau$ with high accuracy. To evaluate the performance of the enhanced PINN model for data-driven parameter discovery in the (2+ 1)-dimensional VC-CNLSE under noisy conditions, we introduce different noise intensities to the observational data. For a fair comparison, we use the same initialization for all networks trained with different noise intensities. As reported in Table 6, the enhanced PINN model accurately predicts the true value of the constant coefficient $\tau$ even when the observational data is corrupted with noise. It is evident that increasing noise levels leads to higher prediction errors and negatively impacts loss convergence. The left panel of Fig. 17 depicts the process of our PINN model learning the true value of the constant coefficient $\tau$. During the iterations of the Adam optimizer (solid lines), followed by the L-BFGS optimizer (dashed lines), our PINN model gradually predicts the true value of $\tau$ with high accuracy. One can see that increasing the noise levels will have a negative impact on the predictions. Training under different noise intensities alters the loss values, as shown in the middle panel of Fig. 17. The $\mathbb{L}^{2}$ relative errors between the constant coefficient $\tau$ and its prediction $\hat{\tau}$ for different levels of added noise are illustrated in the right panel of Fig. 17. It can be seen that as the noise level increases from 0 to $15$, the relative error also increases, and continuing the optimization with the L-BFGS optimizer does not have a significant effect on prediction accuracy. In Tabel 7, we investigate the effect of the number of observations on the prediction accuracy through the lens of $\mathbb{L}^{2}$ relative error. It can be seen that increasing the number of observations reduces the relative error, even in the presence of noisy data. In summary, our enhanced PINN model achieves lower prediction error and faster convergence with clean data (noise $0\%$). However, introducing noise into the data negatively impacts the model’s prediction accuracy and convergence.
@@
Noise ($\%$) Predicted values of $\tau$ $\mathbb{L}^{2}$ relative errors ($\%$) Loss values
0 $0.9998723$ $0.013$ 8.536e038.536e-03
55 $0.9984364$ $0.16$ 1.934e021.934e-02
$10$ $0.98912466$ $1.09$ 4.896e024.896e-02
$15$ $1.01243$ $1.24$ 1.042e011.042e-01
@@
$\bullet$ Data-driven variable parameter discovery
@@
Since these coefficients are functions of the temporal variable $t$, the corresponding neural network should process the temporal coordinate as input data to learn the dynamics of these coefficients. We divide the computational domain $\Omega\times[-2,2]$ into $256\times 256\times 200$ discrete equidistant points and randomly sample $N_{o}=2,000$ data observations from the generated data points. We also sample $N_{r}=20,000$ residual points using the LHS strategy within the computational domain. Additionally, we divide the temporal domain [2,2][-2,2] into $1,000$ equidistant points to assess the accuracy of the underlying PINN model in predicting the variable coefficients through the lens of the $\mathbb{L}^{2}$ relative error. Our PINN model employs two neural networks: the first, parameterized by $\boldsymbol{\theta}_{1}$, takes spatial and temporal coordinates as inputs to train and predict the solutions $\psi_{1}$ and $\psi_{2}$ for forward problem solving; the second, parameterized by $\boldsymbol{\theta}_{2}$, uses only the temporal coordinate as input to predict the variable coefficients. Once these coefficients are predicted, they are directly integrated into the physics-informed components described in (8). Both networks may or may not have the same structure and will be trained simultaneously. In this experiment, the first network has $5$ hidden layers with $40$ neurons in each, while the second network has $3$ hidden layers with $30$ neurons per layer. Then, the underlying PINN model aims to learn the dynamic behavior of both the solutions $\psi_{1}$ and $\psi_{2}$, as well as the dynamic behavior of the variable coefficients, through a minimization problem by minimizing the following loss function: